OpenStudy (anonymous):

how do you go from (2y+1)^2 to (y+0.5)^2 ?

OpenStudy (anonymous):

im taking it someone confused you>

OpenStudy (lgbasallote):

do you mean like proving?? i dont think it can...

OpenStudy (lgbasallote):

if it's 2(y + 0.5)^2 it can...but why do that??

OpenStudy (anonymous):

you are right igbasallote cuz they are not logically equivalent

OpenStudy (anonymous):

The only way that could've happened would be: \[\huge 2(y+0.5)^2.\] the two still would be there, though. This eqn is the same as the other one, though.

OpenStudy (anonymous):

actually just complete the square for the stuff on the left of the equation. i think something canceled with the other stuff in the question lol

OpenStudy (lgbasallote):

maybe we can re-solve it..mind posting it?

OpenStudy (anonymous):

k yeah. the two canceled. study23 how did you do that

OpenStudy (anonymous):

Well, the original equation was: \[\huge (2y+1)^2. \] If you factor out the two, then it is \[\huge 2(y+0.5)^2. \] If you multiply the 2 out, then you get the same answer: \[\huge (2y+1)^2. \] Does that make sense?

OpenStudy (anonymous):

well the question is: find the integral of (4y)/[(2y+1)^2)(y^2+y+1)] dy and i'm just reviewing for my exam but want a quick refresher on the basic question i just asked

OpenStudy (anonymous):

@malcolm11235 I can't help you with calculus - Im only in PreCalc, but I hope that made sense to you...

OpenStudy (anonymous):

yeah except the way i'd go about it is expanding it, factoring out 4, and end up with y^2 + y +0.25. how do i complete the square on that?

OpenStudy (anonymous):

\[(2y+1)^{2}=[2(y+0.5)]^{2}=4(y+0.5)^2\]

OpenStudy (anonymous):

The two expressions are proportional.